# Evaluating the Usability of Microgestures for Text Editing Tasks in Virtual Reality

### v1

@@ Abstract
As virtual reality (VR) continues to evolve, traditional input methods such as handheld controllers and gesture systems often face challenges with precision, social accessibility, and user fatigue. We introduce microGEXT, a lightweight microgesture-based system designed for text editing in VR without external sensors, which utilizes small, subtle hand movements to reduce physical strain compared to standard gestures. We evaluated microGEXT in three user studies. In Study 1 ($N=20$), microGEXT reduced overall edit time and fatigue compared to a baseline system. Study 2 ($N=20$) found that microGEXT performed well in short text selection tasks but was slower for longer text ranges. In Study 3 ($N=10$), participants found microGEXT intuitive for open-ended information-gathering tasks. Across all studies, microGEXT demonstrated enhanced user experience and reduced physical effort, offering a promising alternative to traditional VR text editing techniques.
@@ I Introduction
The gorilla arm effect, common during VR text input, arises from extended arm positions used for virtual keyboard interaction, increasing torque on the shoulder and elbow joints—3.77 times more at the shoulder and twice as much at the elbow compared to relaxed arm positions [3]. Such postures cause fatigue, making tasks like character selection or text editing cumbersome. While large-scale gestures exacerbate this issue, microgestures offer a physically less demanding alternative while retaining the benefits of gesture-based interaction [14, 15]. These subtle, minimal movements reduce physical strain and are well-suited for tasks requiring fine control.
To address these challenges, we introduce microGEXT, a microgesture-based system for text editing in VR. Unlike text entry, text editing benefits more from shortcuts [16] that enable efficient structured interactions. Our microGEXT system leverages built-in VR cameras to detect small, ergonomic movements without external hardware, enabling precise control for tasks such as caret navigation and text selection [17] while minimizing fatigue.
@@ II-A Text Editing in Real World and Virtual Environments
Text selection and editing in immersive virtual environments (VEs) present unique challenges due to the spatial constraints and limitations of traditional input devices like keyboards and mice. Song et al. explored these challenges, highlighting the inefficiencies of conventional devices in VEs and proposing controller-based systems as an alternative to improve text selection performance [18]. De Rosa et al.’s Arrow2edit further addressed the spatial interaction difficulties inherent in VEs, focusing on precision-driven tools for text editing [19].
Directional motion-based text selection systems have also gained attention. Xu et al. developed DMove, a system leveraging directional gestures for more efficient text selection in virtual spaces [8]. Wambecke et al. proposed M[eye]cro, integrating eye-tracking with microgestures to enhance user performance in text selection tasks [20]. Similarly, Dudley et al. introduced the VISAR Keyboard, which employs spatial gestures to augment text selection capabilities [21]. Hajika et al. expanded the toolkit for text selection with RadarHand, a radar-based gesture tracking system designed for high precision in immersive settings [22]. Kim et al. added tactile feedback to the interaction repertoire with VibAware, a system providing physical sensations to support intuitive and responsive text selection [23].
@@
Gestures have emerged as a promising alternative to conventional input devices, particularly in scenarios where traditional methods are impractical. Sellier et al. demonstrated the natural interaction advantages of gesture interfaces for digital manipulation [24], while Wobbrock et al. highlighted the value of user-defined gestures for tailoring interactions to individual preferences, especially on touch surfaces [25]. Le et al. advanced gesture-based workflows with shortcut gestures for text editing, improving efficiency [16]. Combining modalities, Slambekova et al. found that eye-hand gesture integration enhanced text editing speed and accuracy [26]. Beyond optical tracking, Nandakumar et al. proposed FingerIO, a sonar-based gesture tracking system enabling versatile real-world applications [27].
Microgestures have shown significant potential in immersive environments. Chan et al. studied user preferences for microgestures [14], while Faisandaz et al. demonstrated the effectiveness of eyes-free microgestures in immersive contexts [28]. Sharma et al. introduced grasping microgestures for precise text manipulation [29] and SoloFinger for one-handed text editing on mobile devices [30]. Kandoi et al. emphasized intentional microgestures for tasks requiring high precision [31]. Additional studies have explored novel applications of microgestures. Freeman et al. examined rhythmic microgestures for subtle interactions [32], while Soliman et al.’s FingerInput enabled single-hand microgestures in constrained spaces [33]. Tan et al. developed BikeGesture, which extended gesture control to mobile contexts such as biking [34]. Vatavu et al. introduced iFAD, a multitasking framework utilizing microgestures for real-time editing and navigation [35].
@@ III-A Gesture Dataset
We used the Meta Quest Pro VR headset with the XR Hand package [^1] https://docs.unity3d.com/Packages/com.unity.xr.hands@1.4/manual/ to capture hand skeleton data for gesture recognition. In a user-elicitation study on single-hand microgestures, Chan et al. presented a resulting gesture set and identified prevalent conceptual themes among the elicited gestures. Following this dataset, we included seven distinct gestures along with a Null class for a set of text editing tasks in VR (see section III-D).
Ten participants were recruited to interact with a text-editing application, contextualizing each gesture. Before data collection, participants viewed a demonstration video and then performed each gesture 20 times. For static gestures, data was clipped to a standard 2-second duration, while dynamic gestures were clipped to capture complete movement, averaging 5 seconds per gesture. Data was recorded at the Quest Pro’s native frame rate of 72 Hz. We also applied Multi-State Gestures for the dynamic Swipe gesture. This gesture was segmented into sub-states (0–3), with each frame in a clip of length T receiving a corresponding sub-state label, enabling more granular gesture phase analysis.
Participants synchronized their swipe with a visual clock in VR, sliding their index finger from 0% to 100% and back within 5 seconds, allowing for automatic sub-state labeling. Smaller sub-state labels denoted movements near the IndexTip, while larger labels indicated positions closer to the finger base (IndexDistal). Other gestures were assigned a single sub-state label ‘4.’ This sub-state design improved classification and analysis of gesture phases by enabling finer segmentation of dynamic gestures. Also, the ability to ignore Null gestures is important in preventing false activations when the user is not performing intentional gestures since the gesture recognition system is dlimiter-free. We included a Null gesture class to capture common hand movements (e.g., resting hands or pinching) that the user may perform unintentionally between deliberate gestures. Participants were asked to perform gestures while ray casting or pinching, and the data was collected, each lasting approximately 2 seconds.
@@ III-D Gestures and Features
The microGEXT system utilizes a set of microgestures designed to facilitate intuitive and efficient text editing in virtual environments. These gestures are performed with the dominant hand, while the non-dominant hand controls range selection modes [39]. The following sections detail the interaction flow of our selected microgestures and their corresponding features that we chose as suggested by a user elicitation study [14].
@@ III-D1 Cut (Scissor Gesture)
The Scissor gesture mimics a cutting motion with the index and middle fingers, activating the Cut function to remove selected text and store it in the clipboard. Users select text, perform the gesture, and the text is cut—providing an intuitive way to delete content.
@@ III-D3 Caret Navigation and Range Selection (Swipe Gesture)
Inspired by DigitSpace [40] and PinchWatch [41], the Swipe gesture slides the thumb along the index finger, offering:
@@ III-D8 Range Selection Mode (Wrist Rotation)
The left hand controls Range Selection Mode through wrist rotation, cycling between character, word, sentence, and paragraph selection [18]. This intuitive interaction adjusts selection granularity without disrupting workflow [42].
@@ IV-A1 Baseline
In comparison to the microGEXT system, we selected the text editing tool selection method from the OpenXR [^2] https://mbucchia.github.io/OpenXR-Toolkit/ package as the Baseline condition, representing a conventional and well-established approach for tool selection in text editing tasks. This Baseline used a menu-based system for text editing, similar to a pop-up text editing menu on a PC, combined with a pinch gesture to confirm the action. The Baseline was adapted from sample implementations within the OpenXR package. Participants could use ray casting to select the desired text editing tool, such as Copy, Paste, and Cut, and then perform a pinch gesture to confirm the selection. The menu interface design adhered closely to the default OpenXR sample menus, ensuring aesthetic consistency. This Baseline was chosen because it reflects a widely adopted hand interaction technique, allowing users to quickly familiarize themselves with text editing tasks in virtual environments.
@@ IV-A2 Participants and Apparatus
A total of 20 participants (13 male, and 7 female) were recruited from a local university. The age range of the participants was between 17 and 28 years ($M=23.25,SD=2.71$). All participants are students and right-handed. None of the participants involved in data collection participated in the user study. In addition to basic demographic information, all participants reported previous VR experience and the familiarity range was between 1 and 7 ($M=4.25,SD=1.71$). The virtual environment was provided via a Meta Quest Pro VR HMD with hand tracking enabled for interaction. The program was developed in Unity Engine (version 2022.3.3f1c1) with the Open XR Hand package (version 1.4.1). The HMD was connected to a high-performance computer via Quest Link equipped with an Intel i9-13900K CPU, an NVIDIA GeForce RTX 4090 GPU, and 64GB RAM.
@@ IV-B Results
To capture Quantitative Performance, we measured Edit Time for each command and for each round of task, which is the time for each editing attempt, including both successful and reattempted executions; and (2) Reattempts, which is the number of times participants had to re-execute a command, either due to microGEXT misrecognizing the gesture or participants incorrectly performing a gesture or selecting the wrong menu item.
@@ IV-B1 Quantitative Performance
Participants required significantly more reattempts in the microGEXT condition ($M=2.25$, $SD=1.69$) than in the Baseline ($M=0.32$, $SD=0.67$), with a strong effect size ($U=46.5$, $p<0.0001$, $r=-0.657$).
@@ IV-B2 Qualitative Feedback
The UEQ-Short results assessed Pragmatic Quality, Hedonic Quality, and Overall User Experience. For Pragmatic Quality, no significant difference was found between the Baseline ($M=1.4$, $SD=1.12$) and microGEXT ($M=1.35$, $SD=0.99$; $U=212.50$, $p=0.7439$, $r=0.063$). However, for Hedonic Quality, microGEXT significantly outperformed the Baseline ($M=1.96$, $SD=0.70$ vs. $M=-0.05$, $SD=1.47$; $U=40.50$, $p<0.0001$ , $r=0.798$). Similarly, for Overall User Experience, microGEXT ($M=1.66$, $SD=0.70$) was significantly better than the Baseline ($M=0.675$, $SD=1.12$; $U=100.00$, $p=0.0070$, $r=0.500$).
@@ V-A2 Participants and Apparatus
All previous participants ($N=20$) from Study 1 took part in User Study 2 in the same room, following the completion of the final questionnaire in Study 1 and a 5-minute break.
@@ V-B1 Quantitative Performance
In contrast, significant differences were observed for longer text range selection tasks. For the Paragraph task, microGEXT was significantly slower ($M=14.82$ , $SD=4.64$ ) compared to the Baseline ($M=9.96$, $SD=5.67$; $U=62.0$, $p=0.0002$, $r=-0.590$). Similarly, in the Two Paragraphs task, microGEXT ($M=15.51$, $SD=3.29$) was significantly slower than the Baseline ($M=10.38$, $SD=3.94$; $U=40.0$, $p<0.0001$, $r=-0.684$).
@@
When comparing the overall edit time across all tasks, the Baseline condition ($M=17.85$, $SD=11.42$) and microGEXT ($M=16.69$, $SD=3.08$) showed no significant difference ($U=132.0$, $p=0.0679$, $r=-0.291$).
@@ V-C Qualitative Feedback
User Experience Questionnaire
The results from the UEQ-Short for User Study 2 were analyzed using the Mann-Whitney U test. For Pragmatic Quality, there was a significant difference between the Baseline and microGEXT conditions ($U=104.00$, $p=0.0096$, $r=0.4800$). The Baseline condition had a mean score of 0.2625 ($SD=1.55$), while microGEXT had a mean score of 1.5 ($SD=1.04$).
For Hedonic Quality, a highly significant difference was found, with microGEXT outperforming the Baseline ($U=33.5$, $p<0.0001$, $r=0.8325$). The Baseline had a mean of -0.55 ($SD=1.49$), whereas microGEXT had a much higher mean of 1.96 ($SD=0.69$).
In terms of the Overall User Experience, microGEXT also showed a significant improvement over the Baseline ($U=44.00$, $p<0.0001$, $r=0.7800$). The Baseline condition had a mean of -0.14375 ($SD=1.31$), while microGEXT achieved a mean of 1.73 ($SD=0.69$).
@@
The UEQ-Short results for Study 2 revealed significant differences in all dimensions. For Pragmatic Quality, microGEXT ($M=1.5$, $SD=1.04$) scored significantly higher than the Baseline ($M=0.26$, $SD=1.55$; $U=104.00$ , $p=0.0096$, $r=0.480$). Hedonic Quality showed a highly significant improvement with microGEXT ($M=1.96$, $SD=0.69$) outperforming the Baseline ($M=-0.55$, $SD=1.49$; $U=33.50$, $p<0.0001$ , $r=0.833$). For Overall User Experience, microGEXT ($M=1.73$, $SD=0.69$) also significantly surpassed the Baseline ($M=-0.14$, $SD=1.31$; $U=44.00$, $p<0.0001$, $r=0.780$).

### v2

@@ Abstract
As virtual reality (VR) continues to evolve, traditional input methods such as handheld controllers and gesture systems often face challenges with precision, social accessibility, and user fatigue. These limitations motivate the exploration of microgestures, which promise more subtle, ergonomic, and device-free interactions. We introduce microGEXT, a lightweight microgesture-based system designed for text editing in VR without external sensors, which utilizes small, subtle hand movements to reduce physical strain compared to standard gestures. We evaluated microGEXT in three user studies. In Study 1 ($N=20$), microGEXT reduced overall edit time and fatigue compared to a ray-casting + pinch menu baseline, the default text editing approach in commercial VR systems. Study 2 ($N=20$) found that microGEXT performed well in short text selection tasks but was slower for longer text ranges. In Study 3 ($N=10$), participants found microGEXT intuitive for open-ended information-gathering tasks. Across all studies, microGEXT demonstrated enhanced user experience and reduced physical effort, offering a promising alternative to traditional VR text editing techniques.
@@ I Introduction
The gorilla arm effect, common during VR text input, arises from extended arm positions used for virtual keyboard interaction, increasing torque on the shoulder and elbow joints—3.77 times more at the shoulder and twice as much at the elbow compared to relaxed arm positions [3]. Such postures cause fatigue, making tasks like character selection or text editing cumbersome. While large-scale gestures exacerbate this issue, microgestures offer a promising alternative: they are subtle, socially unobtrusive, and physically less demanding, while retaining the benefits of gesture-based interaction [14, 15]. These subtle, minimal movements reduce physical strain and are well-suited for tasks requiring fine control. Recent elicitation studies have also shown that microgestures can be systematically mapped to common text editing commands, providing a principled foundation for our gesture set design [14, 16, 17, 18].
To address these challenges, we introduce microGEXT, a microgesture-based system for text editing in VR. Unlike text entry, text editing benefits more from shortcuts [19] that enable efficient structured interactions. Our microGEXT system leverages built-in VR cameras to detect small, ergonomic movements without external hardware, enabling precise control for tasks such as caret navigation and text selection [20] while minimizing fatigue. For comparison, we evaluate microGEXT against a ray-casting + pinch menu baseline, which is the commercial default method for VR text editing in current head-mounted displays.
@@
Figure 1: In this paper, we introduce the microGEXT system, which replaces traditional menu-based interactions with intuitive microgestures for text editing tasks in virtual environments. This teaser figure highlights the contrast between the conventional text editing menu and the corresponding microgestures used for functions such as mode switching, caret navigation, range selection, cut, copy, paste, select all, undo, and delete (from left to right).
@@ II-A Text Editing in Real World and Virtual Environments
Text selection and editing in immersive virtual environments (VEs) present unique challenges due to the spatial constraints and limitations of traditional input devices like keyboards and mice. Song et al. explored these challenges, highlighting the inefficiencies of conventional devices in VEs and proposing controller-based systems as an alternative to improve text selection performance [21]. De Rosa et al. further addressed the spatial interaction difficulties inherent in VEs, focusing on precision-driven tools for text editing [22].
A broad line of research has therefore explored alternative interaction techniques tailored to immersive contexts. Directional motion-based text selection systems (e.g., DMove [8]) reduce reliance on 2D metaphors by mapping gestures to navigation. Hybrid approaches integrate eye-tracking and gestures, such as M[eye]cro [23], to enhance speed and accuracy. Other systems augment text selection with novel sensing modalities, including spatial gestures [24]), radar-based tracking [25]), and tactile augmentation [26]. Together, these works demonstrate that immersive text editing requires interaction techniques that are precise, ergonomic, and compatible with the physical and social constraints of VR. However, many solutions still rely on large movements or external hardware, which can induce fatigue and limit portability. This gap motivates exploring microgestures as a lightweight and ergonomic alternative.
@@
Gestures have emerged as a promising alternative to conventional input devices, particularly in scenarios where traditional methods are impractical. Sellier et al. demonstrated the natural interaction advantages of gesture interfaces for digital manipulation [27], while Wobbrock et al. highlighted the value of user-defined gestures for tailoring interactions to individual preferences, especially on touch surfaces [28]. For text editing specifically, Le et al. showed that shortcut gestures can significantly improve workflow efficiency [19], and multimodal approaches (e.g., gaze–hand combinations [29]) further enhance precision. Beyond optical tracking, sonar-based systems such as FingerIO [30] illustrate the technical diversity of gesture sensing.
Microgestures extend this line of work by explicitly targeting fatigue reduction and social acceptability in constrained spaces. Chan et al. [14] elicited user-preferred microgestures, highlighting their intuitiveness and minimal effort. Subsequent studies demonstrated effective applications: Faisandaz et al. [31] examined eyes-free use, Sharma et al. explored grasping microgestures [32] and one-handed text editing [33]), and Kandoi et al. [34] emphasized intentionality for high-precision tasks. Additional designs explored rhythmic patterns [35], one-hand use in constrained spaces [36], and mobile contexts such as biking [37]. Frameworks like iFAD further illustrate the potential of microgestures for multitasking [38].
Beyond application prototypes, recent elicitation and systematic studies provide methodological foundations for microgesture design. Chaffangeon Caillet et al. [16] identified transferable microgestures across grasp contexts, Villarreal-Narvaez et al. [17] synthesized 267 elicitation studies into a taxonomy, and Gheran et al. [18] introduced tools for replication and consolidation. These works stress that gesture sets must be justified not only ergonomically but also conceptually, with attention to context and task mappings. Our work builds directly on these insights by designing and evaluating microgestures specifically for VR text editing, a domain that has received limited attention despite its importance for productivity in immersive environments.
@@ III-A Gesture Dataset
We used the Meta Quest Pro VR headset with the XR Hand package [^1] https://docs.unity3d.com/Packages/com.unity.xr.hands@1.4/manual/ to capture hand skeleton data for gesture recognition. Our gesture design was informed by prior elicitation work. In particular, Chan et al. [14] proposed a dataset of user-defined single-hand microgestures, identifying prevalent conceptual themes such as directional swipes, finger touches, and static hand poses. From this set we selected gestures that (i) were rated as comfortable and memorable in prior studies, (ii) mapped naturally to common text editing commands (e.g., Cut, Paste, Select All), and (iii) could be reliably recognized using only the Quest Pro’s built-in cameras. Three gestures (Pinch, Circle, and dynamic Swipe) were newly designed to extend command coverage. For each mapping, we provide a rationale in Section III-D, ensuring semantic alignment with the assigned editing command.
Ten participants were recruited to interact with a text-editing application, contextualizing each gesture. Before data collection, participants viewed a demonstration video and then performed each gesture 20 times. For static gestures, data was clipped to a standard 2-second duration, while dynamic gestures were clipped to capture complete movement, averaging 5 seconds per gesture. Data was recorded at the Quest Pro’s native frame rate of 72 Hz.
To improve recognition, we adopted a multi-state segmentation for the dynamic Swipe gesture. Each clip of length T was divided into sub-states (0–3), automatically labeled based on finger position: smaller values denoted positions near the IndexTip, while larger values indicated the IndexDistal. Other gestures were assigned a single sub-state label ‘4’. This design enabled finer phase-level classification, improving robustness. We also included a dedicated Null class, capturing non-intentional hand movements (e.g., resting pose, incidental pinches). This is critical for delimiter-free recognition, reducing false activations when users are not performing deliberate gestures, as emphasized in prior research [31, 39].
Participants synchronized their swipe with a visual clock in VR, sliding their index finger from 0% to 100% and back within 5 seconds, allowing for automatic sub-state labeling. Data collection lasted approximately 2 seconds per static gesture and 5 seconds for dynamic gestures.
@@ III-D Gestures and Features
Our microGEXT system utilizes a set of microgestures designed to facilitate intuitive and efficient text editing in virtual environments. These gestures are performed with the dominant hand, while the non-dominant hand controls range selection modes [42, 43]. The following sections detail the interaction flow of our selected microgestures and their corresponding features.
@@ III-D1 Cut (Scissor Gesture)
The Scissor gesture mimics a cutting motion with the index and middle fingers, activating the Cut function to remove selected text and store it in the clipboard. Users select text, perform the gesture, and the text is cut.
@@ III-D3 Caret Navigation and Range Selection (Swipe Gesture)
Newly designed for this work, inspired by DigitSpace [44] and PinchWatch [45], the Swipe gesture slides the thumb along the index finger, offering:
@@ III-D8 Range Selection Mode (Wrist Rotation)
Inspired by prior VR 3D modeling work [21] and efficient special character entry [46], the left hand controls Range Selection Mode through wrist rotation, cycling between character, word, sentence, and paragraph selection. This intuitive interaction adjusts selection granularity without disrupting workflow.
@@ IV-A1 Baseline
In comparison to the microGEXT system, we selected the text editing tool selection method from the OpenXR [^2] https://mbucchia.github.io/OpenXR-Toolkit/ package as the Baseline condition, representing a conventional and well-established approach for tool selection in text editing tasks. This Baseline used a ray-casting + pinch menu system, reflecting the default design in many commercial VR applications (e.g., Oculus Browser, Immersed). Participants pointed with ray casting to select the desired tool (e.g., Copy, Paste, Cut) and then confirmed with a pinch gesture. We chose this Baseline because it is the most common default for text editing and command activation in commercial VR systems, thereby providing a familiar and practical comparator rather than a novel or arbitrary technique. The menu interface design adhered closely to the default OpenXR sample menus, ensuring consistency.
@@ IV-A2 Participants and Apparatus
A total of 20 participants (13 male, 7 female) were recruited from a local university. The age range of the participants was between 17 and 28 years ($M=23.25,SD=2.71$). All participants were students, right-handed, and reported previous VR experience (familiarity 1–7, $M=4.25,SD=1.71$). None of the participants involved in data collection participated in the user study. The virtual environment was provided via a Meta Quest Pro VR HMD with hand tracking enabled for interaction. The program was developed in Unity Engine (version 2022.3.3f1c1) with the OpenXR Hand package (version 1.4.1). The HMD was connected to a high-performance computer via Quest Link, equipped with an Intel i9-13900K CPU, an NVIDIA GeForce RTX 4090 GPU, and 64GB RAM. The study protocol was approved by the Cambridge University Engineering Department Ethics Committee (#452), and all participants provided informed consent.
@@ IV-B Results
To capture Quantitative Performance, we measured (1) Edit Time, defined as the duration from when an editing command appeared to its successful execution. This measure included both initial and reattempted executions and was logged automatically by the system; and (2) Reattempts, defined as the number of times a command had to be re-executed due to either microGEXT misrecognition, incorrect gesture performance, or erroneous menu selection.
@@ IV-B1 Quantitative Performance
Participants required significantly more reattempts in the microGEXT condition ($M=2.25$, $SD=1.69$) than in the Baseline ($M=0.32$, $SD=0.67$), with a strong effect size ($U=46.5$, $p<0.0001$, r=0.657r=-0.657).
@@ IV-B2 Qualitative Feedback
The UEQ-Short results assessed Pragmatic Quality, Hedonic Quality, and Overall User Experience. For Pragmatic Quality, no significant difference was found between the Baseline ($M=1.4$, $SD=1.12$) and microGEXT ($M=1.35$, $SD=0.99$; $U=212.50$, $p=0.7439$, $r=0.063$). However, for Hedonic Quality, microGEXT significantly outperformed the Baseline ($M=1.96$, $SD=0.70$ vs. M=0.05M=-0.05, $SD=1.47$; $U=40.50$, $p<0.0001$ , $r=0.798$). Similarly, for Overall User Experience, microGEXT ($M=1.66$, $SD=0.70$) was significantly better than the Baseline ($M=0.675$, $SD=1.12$; $U=100.00$, $p=0.0070$, $r=0.500$).
@@ V-A2 Participants and Apparatus
All previous participants ($N=20$) from Study 1 took part in User Study 2 in the same room, following the completion of the final questionnaire in Study 1 and a 5-minute break. The study protocol was also approved by the Cambridge University Engineering Department Ethics Committee (#452).
@@ V-A3 Study Design and Tasks
Figure 6: Example tasks used in our study 2: (1) Four characters, (2) Four words, (3) One sentence, (4) Two sentences, (5) One paragraph, (6) Two paragraphs.
@@ V-B1 Quantitative Performance
In contrast, significant differences were observed for longer text range selection tasks. For the Paragraph task, microGEXT was significantly slower ($M=14.82$ , $SD=4.64$ ) compared to the Baseline ($M=9.96$, $SD=5.67$; $U=62.0$, $p=0.0002$, r=0.590r=-0.590). Similarly, in the Two Paragraphs task, microGEXT ($M=15.51$, $SD=3.29$) was significantly slower than the Baseline ($M=10.38$, $SD=3.94$; $U=40.0$, $p<0.0001$, r=0.684r=-0.684).
@@
When comparing the overall edit time across all tasks, the Baseline condition ($M=17.85$, $SD=11.42$) and microGEXT ($M=16.69$, $SD=3.08$) showed no significant difference ($U=132.0$, $p=0.0679$, r=0.291r=-0.291).
@@ V-C Qualitative Feedback
The UEQ-Short results for Study 2 revealed significant differences in all dimensions. For Pragmatic Quality, microGEXT ($M=1.5$, $SD=1.04$) scored significantly higher than the Baseline ($M=0.26$, $SD=1.55$; $U=104.00$ , $p=0.0096$, $r=0.480$). Hedonic Quality showed a highly significant improvement with microGEXT ($M=1.96$, $SD=0.69$) outperforming the Baseline (M=0.55M=-0.55, $SD=1.49$; $U=33.50$, $p<0.0001$ , $r=0.833$). For Overall User Experience, microGEXT ($M=1.73$, $SD=0.69$) also significantly surpassed the Baseline (M=0.14M=-0.14, $SD=1.31$; $U=44.00$, $p<0.0001$, $r=0.780$).